I need some help please
If f(x)=4X^3-x symmetric with respect to the x-axis, the y- axis, the origin or none of these
thanks for your time

Answer by stanbon(48568)

Is f(x)=4X^3-x symmetric with respect to the x-axis, the y- axis, the origin or none of these?
------------
f(x)=4x^3-x
-f(x)= -4x^3+x
f(-x)= -4x^3+x
-f(-x)=4x^3-x
--------------
y-axis symmetry if f(x)=f(-x)
Not the same so no y-axis symmetry
-------------
x-axis symmetry if f(x)=-f(x)
Not the same so no x-axis symmetry
---------------
origin symmetry if f(x)=-f(-x)
Yes, so there is origin symmetry
----------
Cheers,
Stan H.